# A Large Language Model-Enabled Control Architecture for Dynamic Resource Capability Exploration in Multi-Agent Manufacturing Systems

### v1

@@ I Introduction
This paper presents an LLM-enabled control architecture for MASs, enabling dynamic exploration of resource capabilities to adapt to real-time disruptions. The major contribution of this paper include: (1) designing a control architecture that enables dynamic decision-making for unexpected disruptions, (2) developing a context-aware system that evaluates the current state of the environment and performance to make informed control decisions, and (3) introducing a mechanism for dynamically exploring resource capabilities, improving agent coordination for system resilience.
@@ II-A Dynamic Adaptation for Manufacturing System Control
Different control strategies have been designed to enhance adaptability in manufacturing. A centralized control approach is used for resource allocation and scheduling [11, 12], leveraging sensor data for decision-making and rerouting parts during disruptions. However, as system complexity grows, centralized architectures face challenges in learning new information as they depend on predefined capabilities [6]. Distributed control strategies, such as MAS, have been introduced to tackle these challenges, where various agents make local decisions under coordination to fulfill global objectives [3, 4]. Recent MAS architectures integrate Product Agents (PAs) and Resource Agents (RAs) to enhance flexibility [13]. PAs manage product-specific decision-making, while RAs control shop-floor resources like robots and machines [13]. Existing MAS solutions improve adaptability by PAs dynamically exploring RAs for task execution [14], while RAs optimize resource allocation during disruptions [15].
@@ II-B Large Language Models in Manufacturing for Adaptation
While these studies highlight the adaptability of LLMs in manufacturing, they focus on simplified processes such as translating basic instructions, process planning, and knowledge retrieval. They do not focus on large-scale manufacturing systems, particularly in multi-agent environments where interactions between agents affect system performance. A gap exists in utilizing LLMs for context-aware decision-making and exploring resource capabilities in response to unexpected disruptions. For instance, if multiple machines operate with varying efficiency levels, LLM should not only retrieve performance data but also dynamically assess current states and constraints to enable reconfigurations. Therefore, we introduce a control architecture that utilizes LLMs for large-scale manufacturing operations, explore capabilities, and enable dynamic decision-making.
@@ III Overview and Problem Statement
The section provides an overview of MAS, which utilizes PAs and RAs, defined in [13, 14] and outlines the resource capability exploration problem.
@@ III-A2 Product Agents
The PA utilizes three main components to support decision-making and interaction with the RA: the process plan, product history, and environment model.
@@
The environment model $M_{e}$ provides PA with a holistic view of the manufacturing environment. The environment model $M_{e}$ is defined as:
$M_{e}=(X_{e},E_{e},Tr_{e},Prp_{p,e},Ag_{h,e},x_{c})$ where $x_{c}$ is the current state of the physical part. $Ag_{h,e}$ is the event-agent association function, mapping each event to the responsible RA. The state set $X_{e}$, event set $E_{e}$, state transition function $Tr_{e}$, and physical properties $Prp_{p,e}$ follow the same definitions in the capabilities model $M$.
Additional details and examples of RA and PA models are provided in [13].
@@ IV-B1 Inputs
LLMs enhance reasoning through rationale exploration, generating and evaluating multiple reasoning paths to improve problem-solving [23]. To ensure structured and reliable outputs, we employ constraint-based prompts [24] and structured Chain-of-Thought (CoT) prompts [25] to enhance the transparency and coherence of LLM decisions. Structured CoT prompts guide the LLM to explicitly outline reasoning steps, clarifying why certain RA or capabilities are selected during disruptions. The decision-maker component utilizes three types of structured inputs for LLM:
@@
  • Template: The template structures the LLM’s output to ensure consistency so that the capabilities can be systematically processed and applied to exploration agent $RA_{e}$.
@@ IV-B2 LLM for Resource Capability Exploration
Various studies have applied “LLM-as-a-judge” across diverse tasks and applications for scoring, ranking, and selection [26]. Although rule-based methods (e.g. selecting the nearest available robot) provide consistent and explainable decisions, they lack adaptability to unforeseen disruptions. Additionally, LLM-based decision-making strategies, such as unconstrained reasoning, lack systematic methods to ensure consistency and reproducibility, which are critical in industrial applications. In this work, we leverage LLMs to evaluate RAs for capability exploration using a weighted scoring function:
@@
where score $S_{i}$ represents the suitability of a candidate resource agent $i$ for exploration based on multiple performance indicators. $\omega_{j}$ is the weight assigned to factor $j$ by the LLM, and $x_{ij}$ is the raw value, such as throughput, utilization, or breakdown rates. The function $f_{j}$ normalizes these values. The agent with the highest score is selected as exploration agent $RA_{e}$, and its capabilities $C_{e}$ are adjusted by exploring the disrupted agent’s functional responsibilities. The scoring approach provides structured, quantifiable comparisons among candidates, aiding transparency and validation. Future work will investigate refinements to prompt design and methodologies to improve decision reliability.
@@ IV-B3 Outputs
Although LLMs are effective for analyzing and generating information, their outputs can be inconsistent. Several studies, such as Self-Refine [27], have employed iterative refinement to reduce errors and enhance consistency. We apply a similar approach, using feedback-driven adjustments to ensure LLM decisions remain within predefined operational and safety constraints. This method is critical in manufacturing contexts, where safety constraints and operational parameters are essential to maintain system stability and reliability.
The validation process, as demonstrated in Algorithm 1, ensures that the selected agent and capabilities comply with both system constraints and safety limits. The validation begins with the CCA retrieving the operational boundary set $\mathcal{B}_{i}$ and safety constraint set $\mathcal{S}_{i}$. Before proceeding, the generated output $O$ is checked for syntax error, ensuring the output $O$ is generated based on the expected structure. If an error is detected, feedback is generated and used to refine the next iteration. The CCA then initializes the validation loop, setting the iteration counter to zero and marking the output as invalid. For each iteration, the CCA extracts information of exploration agent $RA_{e}$ from $O=(RA_{e},C_{e})$. If $RA_{e}$ is not found in the system model, the validation process continues with updated feedback. The corresponding explored capabilities, $C_{e}$ of $RA_{e}$ are then retrieved and compared against $\mathcal{B}_{i}$ and $\mathcal{S}_{i}$. If these conditions hold, the validation is successful, and the output is applied to the system. Otherwise, feedback is provided, and the process continues until a valid response is obtained or the iteration limit $n$ is reached.
@@ IV-D Exploration Process Workflow
The Data Collection stage begins with CCA detecting a disruption and sends a request for handling the event to $RA_{d}$ (Step 1a). The CCA gathers real-time information on the disruption and the resource’s performance (Step 1b). This step is performed for all affected agents within a loop.
@@ IV-E Exploration Example
Figure 5 illustrates an example of dynamic resource capability exploration in MAS with three robots (Robot 1, Robot 2, Robot 3), two machines (Machine 1, Machine 2), and three buffers (Inventory, Product A Buffer, Product B Buffer). A disruption occurs at Robot 2, preventing it from transferring Product A to Machine 1, affecting states $X3$ (at Product A Buffer) and $X4$ (moving to Machine 1). The CCA collects RAs status and performance $PM_{i}(t)$ by retrieving data from RA through direct queries. Metrics, including availability ($A_{i}$), utilization ($U_{i}$), and proximity ($P_{i}$), are stored in a knowledge base for LLM to evaluate potential exploration agents $RA_{e}$. The LLM calculates a score $S_{i}$ using a weighted function $S_{i}=w_{1}A_{i}+w_{2}(1-P_{i})+w_{3}U_{i}$ to select Robot 1 as the exploration agent $RA_{e}$ over Robot 3, based on its higher score. The CCA then validates operational boundaries by checking movement range constraints, confirming that Robot 1 can execute the required operations. The CCA assigns Robot 1 as the $RA_{e}$ and updates its capability model by integrating the disrupted events $\sigma_{3}$ and $\sigma_{4}$. The event set $E^{\prime}$ is modified, and shared states $X_{s}$ are updated, enabling Robot 1 to transition through $X8$ (Inventory to Machine 1) and reach $X5$ (at Machine 1). Through the bidding process, the PA identifies Robot 1 and state $X8$ as the most efficient path to achieve the marked state $X7$ (processed at Machine 1).
@@ V-C Insights from Case Study
The case study highlights that the proposed CCA architecture enhances system resilience and adaptability. This research demonstrated how the system can maintain efficient operations despite disruptions, minimizing part removals and reducing failures. Even when agents experience breakdowns. the system sustains high throughput and minimizes bottlenecks. These findings suggest that the CCA architecture offers a robust and flexible approach for handling dynamic challenges.
@@
Another challenge is the generalizability of the proposed architecture. Although the system successfully handled unexpected robot breakdowns, further evaluation is needed to assess its applicability to other uncertainties, such as machine failures, demand shifts, and simultaneous disruptions of multiple agents. Although fine-tuning LLM can enhance accuracy in manufacturing [18] for specific scenarios, developing an approach capable of handling any disturbances is important for maintaining efficiency. Therefore, the architecture should utilize a knowledge base that is structured for MAS runtime data [30]. This can ensure the system retrieves relevant context related to real-time disturbances with LLM using Retrieval Augmented Generation (RAG) [31].
@@ VI Conclusion and Future Work
This paper introduced a CCA architecture, achieving real-time adaptability (RCh1), enabling context-aware decision-making based on real-time data (RCh2), and relaxing pre-defined operational boundaries by dynamically exploring resource capabilities (RCh3). The proposed framework enhanced real-time adaptation and resource flexibility using LLM. The case study demonstrated that the CCA architecture effectively mitigates disruptions, improving system resilience and operational flexibility. The case study results showed the architecture enabled the system to reassign tasks to exploration agents, minimizing the impact of disruptions on production flow and throughput.
Future work will focus on enhancing the transparency and interpretability of LLM-based decisions to improve system validation and trust in industrial settings. Additional studies will also involve evaluating the architecture against more varied and complex disruptions to evaluate resilience and effectiveness compared to conventional rule-based adaptation methods. The goal is to demonstrate reliable and efficient performance in highly dynamic manufacturing scenarios.

### v2

@@ I Introduction
This paper presents an LLM-enabled control architecture for MASs, enabling dynamic exploration of resource capabilities to adapt to real-time disruptions. The major contribution of this paper include: (1) a control architecture that enables dynamic decision-making for unexpected disruptions, (2) a context-aware system that evaluates the current state of the environment and performance to make informed control decisions, and (3) a novel mechanism for dynamically exploring resource capabilities, improving agent coordination for system resilience.
@@ II-A Dynamic Adaptation for Manufacturing System Control
Different control strategies have been designed to enhance adaptability in manufacturing. A centralized control approach is used for resource allocation and scheduling [11, 12], leveraging sensor data for decision-making and rerouting parts during disruptions. However, as system complexity grows, centralized architectures face challenges in learning new information as they depend on predefined capabilities [6]. Distributed control strategies, such as MAS, have been introduced to address these challenges, where various agents make local decisions under coordination to fulfill global objectives [3, 4]. Recent MAS architectures integrate Product Agents (PAs) and Resource Agents (RAs) to enhance flexibility [13]. PAs manage product-specific decision-making, while RAs control shop-floor resources like robots and machines [13]. Existing MAS solutions improve adaptability by PAs dynamically exploring RAs for task execution [14], while RAs optimize resource allocation during disruptions [15].
@@ II-B Large Language Models in Manufacturing for Adaptation
While these studies highlight the adaptability of LLMs in manufacturing, they focus on simplified processes such as translating basic instructions, process planning, and knowledge retrieval. They do not focus on large-scale manufacturing systems, particularly in multi-agent environments where interactions between agents affect system performance. A gap exists in utilizing LLMs for context-aware decision-making and exploring resource capabilities in response to unexpected disruptions. For instance, if multiple machines operate with varying efficiency levels, LLM should not only retrieve performance data but also dynamically assess current states and constraints to enable reconfigurations. Therefore, we introduce a control architecture that utilizes LLMs to explore capabilities and enable dynamic decision-making in large-scale manufacturing operations.
@@ III Overview and Problem Statement
This section provides an overview of MAS, which utilizes PAs and RAs, and outlines the resource capability exploration problem.
@@ III-A2 Product Agents
The PA utilizes three main components to support decision-making and interaction with the RA: the process plan, product history, and environment model. Additional details and examples of RA and PA models are provided in [13].
@@
The environment model $M_{e}$ provides PA with a holistic view of the manufacturing environment. The environment model $M_{e}$ is defined as: $M_{e}=(X_{e},E_{e},Tr_{e},Prp_{p,e},Ag_{h,e},x_{c})$, where $x_{c}$ is the current state of the physical part. $Ag_{h,e}$ is the event-agent association function, mapping each event to the responsible RA. The state set $X_{e}$, event set $E_{e}$, state transition function $Tr_{e}$, and physical properties $Prp_{p,e}$ follow the same definitions in the capabilities model $M$.
@@ IV-B1 Inputs
LLMs enhance reasoning through rationale exploration, generating and evaluating multiple reasoning paths to improve problem-solving [23]. To ensure structured and reliable outputs, we employ constraint-based prompts [24] and structured Chain-of-Thought (CoT) prompts [25] to enhance the transparency of LLM decisions. Structured CoT prompts guide the LLM to explicitly outline reasoning steps, clarifying why certain RA or capabilities are selected during disruptions. The decision-maker component utilizes three types of structured inputs for LLM:
@@
  • Template: The template structures the LLM’s output to ensure consistency so that the capabilities can be applied to exploration agent $RA_{e}$.
@@ IV-B2 LLM for Resource Capability Exploration
Various studies have applied “LLM-as-a-judge” across diverse applications for scoring, ranking, and selection [26]. Although rule-based methods (e.g. selecting the nearest robot) provide consistent decisions, they lack adaptability to unforeseen disruptions. Additionally, LLM-based decision-making strategies, such as unconstrained reasoning, lack systematic methods to ensure reproducibility. In this work, we leverage LLMs to evaluate RAs for capability exploration using a weighted scoring function:
@@
where score $S_{i}$ represents the suitability of a candidate resource agent $i$ for exploration based on multiple performance indicators. $\omega_{j}$ is the weight assigned to factor $j$ by the LLM, and $x_{ij}$ is the raw value, such as throughput, utilization, or breakdown rates. The function $f_{j}$ normalizes these values. The agent with the highest score is selected as exploration agent $RA_{e}$, and its capabilities $C_{e}$ are adjusted by exploring the disrupted agent’s functional responsibilities. The scoring approach provides structured, quantifiable comparisons among candidates, aiding transparency and validation. In this study, the weights $\omega_{j}$ were determined by the LLM. Systematic and prompt design for explicit weight determination will be done in future work to improve decision reliability.
@@ IV-B3 Outputs
Although LLMs are effective for analyzing and generating information, their outputs can be inconsistent. Several studies, such as Self-Refine [27], have employed iterative refinement to reduce errors. We apply a similar approach, using feedback-driven adjustments to ensure LLM decisions remain within predefined operational and safety constraints. This method is critical in manufacturing contexts, where safety constraints and operational parameters are essential to maintain reliability.
The validation process, as demonstrated in Algorithm 1, ensures that the selected agent and capabilities comply with both system constraints and safety limits. The validation begins with the CCA retrieving the operational boundary set $\mathcal{B}_{i}$ and safety constraint set $\mathcal{S}_{i}$. Before proceeding, the generated output $O$ is checked for syntax error, ensuring the output $O$ is generated based on the expected structure. If an error is detected, feedback is generated and used to refine the next iteration. The CCA then initializes the validation loop, setting the iteration counter to zero and marking the output as invalid. For each iteration, the CCA extracts information of exploration agent $RA_{e}$ from $O=(RA_{e},C_{e})$. If $RA_{e}$ is not found, the validation process continues with updated feedback. The corresponding explored capabilities, $C_{e}$ of $RA_{e}$ are then retrieved and compared against $\mathcal{B}_{i}$ and $\mathcal{S}_{i}$. If these conditions hold, the validation is successful, and the output is applied. Otherwise, feedback is provided, and the process continues until a valid response is obtained or the iteration limit $n$ is reached.
@@ IV-D Exploration Process Workflow
The Data Collection stage begins with CCA detecting a disruption and sends a request for handling the event to $RA_{d}$ (Step 1a). The CCA gathers real-time information on the disruption and the performance (Step 1b). This step is performed for all affected agents within a loop.
@@ IV-E Exploration Example
Figure 5 illustrates an example with three robots (Robot 1, Robot 2, Robot 3), two machines (Machine 1, Machine 2), and three buffers (Inventory, Product A Buffer, Product B Buffer). A disruption occurs at Robot 2, preventing it from transferring Product A to Machine 1, affecting states $X3$ (at Product A Buffer) and $X4$ (moving to Machine 1). The CCA collects RAs status and performance $PM_{i}(t)$ by retrieving data from RA through direct queries. Metrics, including availability ($A_{i}$), utilization ($U_{i}$), and proximity ($P_{i}$), are stored in a knowledge base for LLM to evaluate potential exploration agents $RA_{e}$. For example, the user provides the following prompt:
"Analyze real-time data (R1: proximity=2, utilization=.92; R2: proximity=6, utilization=.35...) to select an exploration agent(availability, utilization, proximity). Enforce movement constraints and output your decision with clear reasoning..."
Based on this instruction, the LLM calculates a score $S_{i}$ using a weighted function $S_{i}=w_{1}A_{i}+w_{2}(1-P_{i})+w_{3}U_{i}$ to select Robot 1 as the exploration agent $RA_{e}$ over Robot 3, based on its higher score. The CCA then validates operational boundaries by checking movement range constraints, confirming that Robot 1 can execute the required operations. The CCA assigns Robot 1 as the $RA_{e}$ and updates its capability model by integrating the disrupted events $\sigma_{3}$ and $\sigma_{4}$. The event set $E^{\prime}$ is modified, and shared states $X_{s}$ are updated, enabling Robot 1 to transition through $X8$ (Inventory to Machine 1) and reach $X5$ (at Machine 1). Through the bidding process, the PA identifies Robot 1 and state $X8$ as the most efficient path to achieve the marked state $X7$ (processed at Machine 1).
@@ V-C Insights from Case Study
The case study highlights that the proposed CCA architecture enhances system resilience and adaptability. This research demonstrated how the system can maintain efficient operations despite disruptions, minimize part removals, and reduce failures. Even when agents experience breakdowns, the system sustains stable throughput and minimizes bottlenecks. These findings suggest that the CCA architecture offers a robust and flexible approach for handling dynamic challenges.
@@
Another challenge is the generalizability of the proposed architecture. Although the system successfully handled unexpected robot breakdowns, further evaluation is needed to assess its applicability to other uncertainties, such as machine failures, demand shifts, and simultaneous disruptions of multiple agents. Although fine-tuning LLM can enhance accuracy in manufacturing [18] for specific scenarios, developing an approach capable of handling any disturbances is important for maintaining efficiency. Additionally, the accuracy of LLM responses relies on structured and relevant runtime data. Therefore, the architecture should utilize a knowledge base that is designed for MAS runtime control [30]. This can ensure the system retrieves relevant context related to real-time disturbances with LLM using Retrieval Augmented Generation (RAG) [31].
@@ VI Conclusion and Future Work
This paper introduced a CCA architecture, achieving real-time adaptability (RCh1), enabling context-aware decision-making based on real-time data (RCh2), and relaxing pre-defined operational boundaries by dynamically exploring resource capabilities (RCh3). The proposed framework enhanced real-time adaptation and resource flexibility using LLM. The case study demonstrated that the CCA architecture effectively mitigates disruptions, improving system resilience and operational flexibility. The case study results showed the architecture enabled the system to reassign tasks to exploration agents, significantly reducing downtime costs in high-mix manufacturing environments.
Future work will focus on enhancing the transparency and interpretability of LLM-based decisions to improve system validation and trust in industrial settings. Additional studies will involve evaluating the architecture against more varied and complex disruptions in real-world manufacturing scenarios to assess resilience and effectiveness compared to conventional rule-based adaptation methods. To further ensure the practicality of the proposed system, future research will explore the impact of LLM response latency in real-world manufacturing environments. The goal is to demonstrate dependable and efficient performance in highly dynamic and practical manufacturing scenarios.